Solve the formula:
P=A/1+ar solve for r
Solve the equation:
1/x-2 - 1/x+2 = 1/15

Answer by vleith(2983)

P=A/1+ar solve for r
Make sure that this is NOT P = A/(l+ar).

P=A/1+ar
P - A/l = ar
(P-A/l)/a = r


Problem 2
1/x-2 - 1/x+2 = 1/15
((x+2) - (x-2))/(x-2)(x+2) = 1/15
(x+2-x+2)* 15 = (x-2)(x+2)
60 = x^2 -4
64 = x^2
x = 8 or -8